60. solve the equation algabraically log.4 x-log.4(x-1)=1/2
Answer by vleith(2983)

log.4 x-log.4(x-1)=1/2
log.4(x/(x-1)) = 1/2
log(x/(x-1))/log(4) = 1/2
log(x/(x-1))/0.602 = 1/2
log(x/(x-1)) = 0.301
x/(x-1) = 10^(0.0301)
x/(x-1) = 2
x = 2x -2
x = 2

Try that out in the original equation, using base transformation
log(2)/log(4) - log(1)/log(4) = 1/2